At 9:00 A.M., your friends begin hiking at 2 mi/h. You begin from the same place at 9:25 A.M. You hike at 3 mi/h.
a. How long will you hsve hiked when you catch up with your friends?
b. At what time will you catch up with your friends?

Answer by Alan3354(69443)

At 9:00 A.M., your friends begin hiking at 2 mi/h. You begin from the same place at 9:25 A.M. You hike at 3 mi/h.
a. How long will you hsve hiked when you catch up with your friends?
In 25 mins, they've gone 2 mi/hr * (25/60) hr = 5/6 miles
You gain on them at 1 mph (3 - 2) so it'll take you 5/6 hours to catch them, or 50 minutes.
5/6 hours * 3 mph = 5/2 miles = 2.5 miles
---------------------------------------
b. At what time will you catch up with your friends?
0925 + 50 mins = 1015